OpenStudy (lord2clash):

Which point lies on the line described by the equation below? y + 3 = 2(x - 1) A. (1, -3) B. (0, 0) C. (2, 1) D. (2, 9) E. (1, -4) F. (-1, -7)

pooja195 (pooja195):

\[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)

pooja195 (pooja195):

We know that y1 is the y in a coordinate which number is it?

OpenStudy (lord2clash):

3

pooja195 (pooja195):

Good but remember that 2 negatives make a positive therefore it would be -3

pooja195 (pooja195):

That should give you the answer only on of your answer choices has -3 in it :P

OpenStudy (lord2clash):

thank you cx

pooja195 (pooja195):

yw ^-^